OpenStudy (anonymous):

pleaseeee? When 36 grams of aluminum carbide react with an excess of water, as shown in the unbalanced chemical equation below, how many moles of aluminum hydroxide will be produced? Al4C3 + H2O Al(OH)3 + CH4 0.25 moles 0.50 moles 1.0 moles 2.0 moles

OpenStudy (anonymous):

first find the molar mass of aluminum carbide. Al * 4 + C * 3. then take the mass divide by the molar mass. be sure to check the unit

OpenStudy (anonymous):

i got 35.9

OpenStudy (anonymous):

ok first lets start with balancing reaction equation: Al4C3 + 12 H2O --> 4 Al(OH)3 + 3 CH4 now we can start to calculate, and first step is to calculate the given amount of Al4C3 from grams in moles: n= m/M n(Al4C3) = m(Al4C3)/M(Al4C3) = 36 g / 143,953 g mol-1 n(Al4C3) = 0,25 mol now we need to take a look at balanced reaction and find out the ratio: n(Al(OH)3)/n(Al4C3) = 1/4 n(Al(OH)3)= 4/1 * n(Al4C3) = 4 * 0,25mol = 1,0 mol